OpenStudy (anonymous):

find the volume B bounded by the paraboloid z=4-x^2-y^2 and the xy plane

OpenStudy (dumbcow):

 \[V = \pi \int\limits_0^4 r^2 = \pi \int\limits_0^4 (4-z) dz\]